# Finite-Time Behavior of Erlang-C Model: Mixing Time, Mean Queue Length and Tail Bounds

### v3

@@ 2.3. Main Results
Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system with the arrival rate $\lambda_{n}=n-n^{1-\alpha}$ and service rate $1$ and let the stationary distribution be $\nu_{n}$. For $\alpha\geq 1$, we have that:
(6) $\chi(\pi_{n,t},\nu_{n})\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0.$
For $\alpha\in(1/2,1)$ and $n\geq 65$, we have that:
(7) $\chi(\pi_{n,t},\nu_{n})\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0,$
@@
For $\alpha\in(0,1/2)$ and a sufficiently large $n$ such that $\lambda_{n}\geq 3$, we have that:
(8) $\chi(\pi_{n,t},\nu_{n})\leq e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0,$
@@
(9) $\chi(\pi_{n,t},\nu_{n})\leq e^{-\frac{t}{1387444804}}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0.$
The proof of this theorem and the exact expressions of $C_{n},D_{n}$ are described in Appendix B, and its proof sketch is summarized in Section 4. Note that we assume $n\geq 65$ for $\alpha\in[1/2,1)$ only to obtain an explicit mixing rate for $\alpha=1/2$ and to get reasonable constants in the proof. Obtaining an explicit bound for all $n\geq 1$ is possible at the cost of worse universal constants in our mixing rate bounds.
It is well-known that there is a phase transition at $\alpha=1/2$ (Halfin and Whitt, 1981), and the behavior in the regimes $\alpha\in(0,1/2)$ and $\alpha\in(1/2,\infty)$ is similar to $M/M/\infty$ and $M/M/1$ respectively. Our Theorem 3 captures such a phase transition, as evident by the mixing-time bounds for the $M/M/1$ and $M/M/\infty$ queues stated below:
@@
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda>0$ and the service rate $\mu>0$ such that $\lambda/\mu\in\mathbb{Z}^{+}$ and let the stationary distribution be $\nu_{\infty}$, we have:
@@
While the above results for $M/M/1$ and $M/M/\infty$ are previously established in (Chafaï, 2006; Morse, 1955; Robert, 2003), it is worth noting that our Lyapunov-Poincaré methodology provides a unified approach to obtain mixing results for $M/M/1$, $M/M/n$, and $M/M/\infty$ queues. The proof of Proposition 2, presented in Appendix B.1.4, is instructive to understand the case of $\alpha\in(1/2,\infty]$ of Theorem 3. Similarly, the proof of Proposition 3, while it is only applicable for $\lambda\in\mathbb{Z}^{+}$, in Appendix B.5 is helpful to understand the case of $\alpha\in(0,1/2)$. Observe that the mixing rate of the $M/M/n$ queue for $\alpha\in(1/2,\infty]$ and the $M/M/1$ queue is $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}\approx n^{1-2\alpha}\rightarrow 0$ as $n\rightarrow\infty$. On the other hand, the mixing rate of the $M/M/n$ queue for $\alpha\in(0,1/2)$ and the $M/M/\infty$ queue is $\Theta(1)$ as $n\rightarrow\infty$. Thus, our theorem characterizes the phase transition of the mixing rate from 00 to $\Theta(1)$ at $\alpha=1/2$. Finally, consider the case of $\alpha=1/2$, similar to the case of $\alpha\in(0,1/2)$, our mixing rate is $\Theta(1)$ but is off by a constant. For this case, a precise characterization of the mixing rate in the limit as $t\rightarrow\infty$, and $n\rightarrow\infty$ was provided in (Gamarnik and Goldberg, 2013), and obtaining a finite time behavior for a finite-sized system that exactly matches these limiting results is an interesting future direction.
While we establish a $\Theta(1)$ bound on the mixing rate for $\alpha\in(0,1/2)$, it is off by a constant compared to the mixing rate of $M/M/\infty$ queue, indicating that our mixing rate characterization is loose in this regime. Nonetheless, we believe our characterization of the mixing rate is state-of-the-art. For example, the result of (Zeifman, 1991) implies a mixing rate of $1-\frac{\lambda_{n}}{n-1}=O\left(n^{-\alpha}\right)\rightarrow 0$ as $n\rightarrow\infty$. Moreover, by imposing additional technical assumptions, we are able to improve our results to obtain tight mixing rates in this regime. To this end, we present two results assuming that either $\lambda_{n}$ is very small (light traffic, i.e. $\lambda_{n}/n\rightarrow 0$) or $\lambda_{n}\in\mathbb{Z}^{+}$, where the latter is an artifact of our methodology as discussed in Section 3.
@@
for some positive parameter $L_{n}$ such that $\lim_{n\rightarrow\infty}L_{n}=1-\sqrt{c}$.
Since $\lambda_{n}\in(0,n)$, there exists a unique positive number $\alpha_{n}$ such that $\lambda_{n}=n-n^{1-\alpha_{n}}$, and so the assumption $\lim_{n\rightarrow\infty}\lambda_{n}/n=c$ means that $\lim_{n\rightarrow\infty}\alpha_{n}\log n=-\log(1-c)$. When $c=0$, we have the mixing rate of the system matches the limiting behavior at the asymptotic (that is, mixing rate approaches $1$ as $n\rightarrow\infty$). On the other hand, for $c\in(0,1)$, Proposition 4 shows that the system admits a mixing rate that is bounded below by a constant that is independent of $n$. Moreover, when $c<\frac{576}{625}$, we get a better constant than the $\frac{1}{25}$ constant in the $\alpha\in(0,1/2)$ regime. One can find the proof of Proposition 4 in Appendix B.4. However, since we know that the phase transition happens at the Halfin-Whitt regime, this suggests that we could do better by obtaining an analysis that matches the limiting behavior for all $\alpha\in(0,1/2)$, possibly with some additional conditions. And so, we consider the case when the arrival rate is an integer, i.e. there is a sequence $\{\lambda_{k}\}_{k\geq\mathbb{Z}^{+}}$ such that $\lambda_{k}\in\mathbb{Z}\,\forall k\in\mathbb{Z}^{+}$. In this case, we can show that as $\lambda_{k}\rightarrow\infty$, we have the mixing rate approaches $1$ as follows.
@@
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the steady state distribution of the continuous-time $M/M/n$ system with unit service rate respectively and let $\{\lambda_{n}\}$ be a sequence of integer arrival rates ($\lambda_{n}\in\mathbb{Z}$) such that $\lambda_{n}\geq 0,n-\lambda_{n}\geq 1$ and
(10) $\lim_{n\rightarrow\infty}\frac{\log\left(n-\lambda_{n}\right)}{\log n}=1-\alpha$
@@
The proof of Proposition 5 is presented in Appendix B.3.3. With additional technical assumptions on $\lambda$, our mixing rate now matches with that of the $M/M/\infty$ queue asymptotically for $\alpha\in(0,1/2)$. Now, when both of these conditions are not satisfied, we can still show mixing but the mixing rate is only tight up to a constant, as shown in Theorem 3. Nevertheless, we believe that these two conditions are rather artificial and we conjecture that these conditions can be lifted to get a mixing rate approaching $1$, matching the mixing rate of the $M/M/\infty$ at the asymptotics.
2.3.1. Finite-time statistics
@@
(12) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n}).$
@@
(13) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$
@@
(14) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-D_{n}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$
@@
And finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$, then we have
(15) $\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-L_{n}t}n\chi(\pi_{n,0},\nu_{n})$
for some $L_{n}>0$ such that $\lim_{n\rightarrow\infty}L_{n}=1$.
One can prove Corollary 4 by combining the Theorem 3 and Corollary 1 for $k=1$. We defer the details of the proof to Appendix C.1. Similar to the above corollary, we can obtain bounds on any moment $k\in\mathbb{Z}^{+}$ by Corollary 1 combined with a $2k$-moment bound on the steady-state distribution $\nu_{n}$. Moreover, in addition to the mean queue length bound, we can establish some tail-bound results to gain some understanding on the concetration of the finite-time behavior of the queue length distribution as follows.
@@
And finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=0$ then we have
$\mathbb{P}_{\pi_{n,t}}\left[q-n>x\right]$ $\leq\left(1+e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{n^{1-\alpha_{n}}}{2}}+4n^{\alpha_{n}-1}+\frac{8n^{2\alpha_{n}-1}}{e}}\times e^{-\frac{x}{2}}$
where $\lim_{n\rightarrow\infty}\alpha_{n}=0$ and $L_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}L_{n}=1$.
The proof of Corollary 5 is deferred to Appendix C.2. Note that if $q(t)\geq n$, then $q(t)-n$ is the number of waiting customers at time $t$. The results in Corollary 5 tell us what is the tail distribution of the number of waiting customers at time $t$. In the Sub-Halfin-Whitt regime, we have the pre-exponent term approaches 00 as $n\rightarrow\infty$, which suggests a faster decay rate than exponential. Moreover, we can also estimate the probability of having an idle server at time $t$ as follows.
@@
The proof of Corollary 6 can be found in Appendix C.3. These results are the finite-time version of the steady-state results established in (Jhunjhunwala et al., 2023) and thus can be seen as the generalization of such results.
3. The Lyapunov-Poincaré method

### v4

@@ 2.3 Main Results
Let $\pi_{n,t}$ be the queue length distribution at time $t\geq 0$ of the continuous-time $M/M/n$ system defined above and let the stationary distribution be $\nu_{n}$.
For $\alpha\geq 1,n\geq 2$, we have that:
$\chi(\pi_{n,t},\nu_{n})\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0.$ (6)
For $\alpha\in(1/2,1),n\geq 2$, we have that:
$\chi(\pi_{n,t},\nu_{n})\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0,$ (7)
@@
For $\alpha\in(0,1/2),n\geq 2$, we have that:
$\chi(\pi_{n,t},\nu_{n})\leq e^{-D_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0,$ (8)
@@
$\chi(\pi_{n,t},\nu_{n})\leq e^{-H_{n}t}\chi(\pi_{n,0},\nu_{n})\,\forall t\geq 0$ (9)
for some $H_{n}>\frac{1}{10861}$ and $\lim_{n\rightarrow\infty}H_{n}=\frac{1}{1781}$.
The proof of this theorem and the exact expressions of $C_{n},D_{n},H_{n}$ are described in Appendix 9, and its proof sketch is summarized in Section 4. Note that we assume $n\geq 2$ so that $\lambda_{n}>0$ (for the convergence rate of $M/M/1$, please refer to Proposition 2), and assume $n\geq 110$ for $\alpha=1/2$ only to obtain an explicit mixing rate for this regime and to get reasonable constants in the proof. Obtaining an explicit bound for all $n\geq 1$ is possible at the cost of worse universal constants in our mixing rate bounds. Now, denote $\tau_{mix}(\varepsilon)=\min_{t}\chi(\pi_{n,t},\nu_{n})\leq\varepsilon$, we obtain the following corollary on the mixing time of the system from the bound on the Chi-square distance.
Corollary 3.
(Mixing time bounds for $M/M/n$ systems) Let $\pi_{n,t}$ be the queue length distribution at time $t$ of the continuous-time $M/M/n$ system defined above and let the stationary distribution be $\nu_{n}$. For $\alpha\geq 1$, we have that:
$\tau_{mix}(\varepsilon)\leq 4n^{2\alpha-1}\log\left(\frac{\chi(\pi_{n,0},\nu_{n})}{\varepsilon}\right).$ (10)
For $\alpha\in(1/2,1)$, we have that:
$\tau_{mix}(\varepsilon)\leq 4C_{n}^{-1}n^{2\alpha-1}\log\left(\frac{\chi(\pi_{n,0},\nu_{n})}{\varepsilon}\right)$ (11)
for some $C_{n}>0$ such that $\lim_{n\rightarrow\infty}C_{n}=1$.
For $\alpha\in(0,1/2)$, we have that:
$\tau_{mix}(\varepsilon)\leq D_{n}^{-1}\log\left(\frac{\chi(\pi_{n,0},\nu_{n})}{\varepsilon}\right)$ (12)
for some $D_{n}>0$ such that $\lim_{n\rightarrow\infty}D_{n}=1/25$.
Finally, for $\alpha=1/2$ and $n\geq 110$, we have
$\tau_{mix}(\varepsilon)\leq H_{n}^{-1}\log\left(\frac{\chi(\pi_{n,0},\nu_{n})}{\varepsilon}\right)$ (13)
for some $H_{n}>\frac{1}{10861}$ and $\lim_{n\rightarrow\infty}H_{n}=\frac{1}{1781}$.
Corollary 3 tells us that the $\varepsilon$-mixing time $\tau_{mix}(\varepsilon)$ has logarithmic dependence on the initial distance and $\varepsilon^{-1}$. Additionally, when the traffic is light (i.e. $\alpha\in(0,1/2)$), it takes $O(1)$ time to mix as $D_{n},H_{n}=\Omega(1)$. However, when the system has heavy traffic (i.e. $\alpha>1/2$), $n^{2\alpha-1}$ would grow to infinity as $n$ grows, and so, it would take a long time to mix in this regime. We defer the proof of Corollary 3 to Appendix 10. Now, it is well-known that there is a phase transition at $\alpha=1/2$ [32], and the behavior in the regimes $\alpha\in(0,1/2)$ and $\alpha\in(1/2,\infty)$ is similar to $M/M/\infty$ and $M/M/1$ respectively. Our Theorem 1 captures such a phase transition, as evident by the mixing-time bounds for the $M/M/1$ and $M/M/\infty$ queues stated below:
@@
Let $\pi_{t,\infty}$ be the queue length distribution at time $t$ of the continuous-time $M/M/\infty$ system with the arrival rate $\lambda$ and the service rate $\mu$ and let the stationary distribution be $\nu_{\infty}$, we have if $\lambda/\mu\in\mathbb{Z}^{+}$ then:
@@
While the above results for $M/M/1$ and $M/M/\infty$ are previously established in [12, 54, 67], it is worth noting that our Lyapunov-Poincaré methodology provides a unified approach to obtain mixing results for $M/M/1$, $M/M/n$, and $M/M/\infty$ queues. The proof of Proposition 2, presented in Appendix 9.1.4, is instructive to understand the case of $\alpha\in(1/2,\infty]$ of Theorem 1. Similarly, the proof of Proposition 3, while it is only applicable for $\lambda\in\mathbb{Z}^{+}$, in Appendix 9.5 is helpful to understand the case of $\alpha\in(0,1/2)$. Observe that the mixing rate of the $M/M/n$ queue for $\alpha\in(1/2,\infty]$ and the $M/M/1$ queue is $(\sqrt{n}-\sqrt{\lambda_{n}})^{2}\approx n^{1-2\alpha}\rightarrow 0$ as $n\rightarrow\infty$. On the other hand, the mixing rate of the $M/M/n$ queue for $\alpha\in(0,1/2)$ and the $M/M/\infty$ queue is $\Theta(1)$ as $n\rightarrow\infty$. Thus, our theorem characterizes the phase transition of the mixing rate from 0 to $\Theta(1)$ at $\alpha=1/2$. Finally, consider the case of $\alpha=1/2$, similar to the case of $\alpha\in(0,1/2)$, our mixing rate is $\Theta(1)$ but is off by a constant. For this case, a precise characterization of the mixing rate in the limit as $t\rightarrow\infty$, and $n\rightarrow\infty$ was provided in [29], and obtaining a finite time behavior for a finite-sized system that exactly matches these limiting results is an interesting future direction.
While we establish a $\Theta(1)$ bound on the mixing rate for $\alpha\in(0,1/2)$, it is off by a constant compared to the mixing rate of $M/M/\infty$ queue, indicating that our mixing rate characterization is loose in this regime. Nonetheless, we believe our characterization of the mixing rate is state-of-the-art. For example, the result of [86] implies a mixing rate of $1-\frac{\lambda_{n}}{n-1}=O\left(n^{-\alpha}\right)\rightarrow 0$ as $n\rightarrow\infty$. Moreover, by imposing additional technical assumptions, we are able to improve our results to obtain tight mixing rates in this regime. To this end, we present two results assuming that either $\lambda_{n}$ is very small (light traffic, i.e. $\lambda_{n}/n\rightarrow 0$) or $\lambda_{n}\in\mathbb{Z}^{+}$, where the latter is an artifact of our methodology as discussed in Section 3.
@@
for some positive parameter $L_{n}$ such that $\lim_{n\rightarrow\infty}L_{n}=1-c$.
Since $\lambda_{n}\in(0,n)$, there exists a unique positive number $\alpha_{n}$ such that $\lambda_{n}=n-n^{1-\alpha_{n}}$, and so the assumption $\lim_{n\rightarrow\infty}\lambda_{n}/n=c$ means that $\lim_{n\rightarrow\infty}\alpha_{n}\log n=-\log(1-c)$. When $c=0$, we have the mixing rate of the system matches the limiting behavior at the asymptotic (that is, mixing rate approaches $1$ as $n\rightarrow\infty$). On the other hand, for $c\in(0,1)$, Proposition 4 shows that the system admits a mixing rate that is bounded below by a constant that is independent of $n$. Moreover, when $c<\frac{24}{25}$, we get a better constant than the $\frac{1}{25}$ constant in the $\alpha\in(0,1/2)$ regime. One can find the proof of Proposition 4 in Appendix 9.4 and the discussion on this result in comparison with previous works in Appendix 12. However, since we know that the phase transition happens at the Halfin-Whitt regime, this suggests that we could do better by obtaining an analysis that matches the limiting behavior for all $\alpha\in(0,1/2)$, possibly with some additional conditions. And so, we consider the case when the arrival rate is an integer, i.e. there is a sequence $\{\lambda_{k}\}_{k\geq\mathbb{Z}^{+}}$ such that $\lambda_{k}\in\mathbb{Z}\,\forall k\in\mathbb{Z}^{+}$. In this case, we can show that as $\lambda_{k}\rightarrow\infty$, we have the mixing rate approaches $1$ as follows.
@@
Let $\pi_{n,t},\nu_{n}$ be the queue length distribution at time $t$ and the steady state distribution of the continuous-time $M/M/n$ system with unit service rate respectively and let $\{\lambda_{n}\}$ be a sequence of integer arrival rates ($\lambda_{n}\in\mathbb{Z}$) such that $\lambda_{n}\geq 0,n-\lambda_{n}\geq 1,\frac{\log\left(n-\lambda_{n}\right)}{\log n}\in(1/2,1)$ and
$\lim_{n\rightarrow\infty}\frac{\log\left(n-\lambda_{n}\right)}{\log n}=1-\alpha$ (14)
@@
The proof of Proposition 5 is presented in Appendix 9.3.3. With additional technical assumptions on $\lambda$, our mixing rate now matches with that of the $M/M/\infty$ queue asymptotically for $\alpha\in(0,1/2)$. Now, when both of these conditions are not satisfied, we can still show mixing but the mixing rate is only tight up to a constant, as shown in Theorem 1. Nevertheless, we believe that these two conditions are rather artificial and we conjecture that these conditions can be lifted to get a mixing rate approaching $1$, matching the mixing rate of the $M/M/\infty$ at the asymptotics.
2.3.1 Finite-time statistics
@@
$\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n}).$ (16)
@@
$\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-C_{n}(\sqrt{n}-\sqrt{\lambda_{n}})^{2}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$ (17)
@@
$\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-D_{n}t}\sqrt{2}\left(n+n^{\alpha}\right)\chi(\pi_{n,0},\nu_{n})$ (18)
@@
And finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$ for $c\in[0,1)$, then we have
$\left|\mathbb{E}_{\pi_{n,t}}[q]-\mathbb{E}_{\nu_{n}}[q]\right|\leq e^{-L_{n}t}n\chi(\pi_{n,0},\nu_{n})$ (19)
for some $L_{n}>0$ such that $\lim_{n\rightarrow\infty}L_{n}=1-c$.
One can prove Corollary 4 by combining the Theorem 1 and Corollary 1 for $k=1$. We defer the details of the proof to Appendix 11.1. Similar to the above corollary, we can obtain bounds on any moment $k\in\mathbb{Z}^{+}$ by Corollary 1 combined with a $2k$-moment bound on the steady-state distribution $\nu_{n}$. Moreover, in addition to the mean queue length bound, we can establish some tail-bound results to gain some understanding on the concetration of the finite-time behavior of the queue length distribution as follows.
@@
Finally, if $\lambda_{n}$ is a sequence of arrival rates such that $\lim_{n\rightarrow\infty}\frac{\lambda_{n}}{n}=c$ for $c\in[0,1)$ then we have
$\mathbb{P}_{\pi_{n,t}}\left[q-n>x\right]$ $\leq\left(1+e^{-L_{n}t}\chi(\pi_{n,0},\nu_{n})\right)\sqrt{e^{-\frac{(1-c)n^{1-\alpha_{n}}}{2}}+2n^{\alpha_{n}-1}+\frac{8n^{2\alpha_{n}-1}}{e(1-c)}}\times e^{-\frac{x}{2}}$
where $\lim_{n\rightarrow\infty}\alpha_{n}=0$ and $L_{n}$ is a positive parameter such that $\lim_{n\rightarrow\infty}L_{n}=1-c$.
The proof of Corollary 5 is deferred to Appendix 11.2. Note that if $q(t)\geq n$, then q(t)nq(t)-n is the number of waiting customers at time $t$. The results in Corollary 5 tell us what is the tail distribution of the number of waiting customers at time $t$. In the Sub-Halfin-Whitt regime, we have the pre-exponent term approaches 0 as $n\rightarrow\infty$, which suggests a faster decay rate than exponential. Moreover, we can also estimate the probability of having an idle server at time $t$ as follows.
@@
The proof of Corollary 6 can be found in Appendix 11.3. These results are the finite-time version of the steady-state results established in [35] and thus can be seen as the generalization of such results.
3 The Lyapunov-Poincaré method